mikewwe13:

The point (-6, y) lies on the graph of the equation: 5y=2x−4. What is the value of y? A. -1/65 B. −5/16 C. 5/16 D. 16/5

Vocaloid:

(-6,y) means the x-coordinate is -6 so simply plug in x = -6 to get y 5y=2x−4

mikewwe13:

the answer is either A or C

Vocaloid:

plug in x = -6 into the equation and tell me what you get for y there is only one correct answer.

mikewwe13:

The answer is B

mikewwe13:

correct me if i'm wrong

Vocaloid:

x = -6 means replace "x" with "-6" then solve for y.

Vocaloid:

please show the steps of your calculations so I can see where you are making mistakes

mikewwe13:

y = -2

Vocaloid:

that is not one of the answer choices please show all steps of your calculations so I can see where you are going wrong.

mikewwe13:

no it's not i know that but i just solved it for y and my calculations shows the answer is D

Vocaloid:

again, please write out the actual steps of your calculations.

Vocaloid:

I'm not going to judge you, I just want to know where you are making mistakes so I can teach you the right way.

mikewwe13:

but also 5y=2(-6)-4 5y=-12-4 5y=-16 y=-16/5

Vocaloid:

good so the answer is y= -16/5

mikewwe13:

The following graph represents the equation y = 1/5x + 1/5 https://d3upxwx9bggvi1.cloudfront.net/img/questions/arei10_1_007.png Which ordered pair is a solution to the equation? A. (0, 0) B. (4, 1) C. (1, 4) D. (1/5, 1/5)